OpenStudy (anonymous):

7. What is the slope of the line that passes through the pair of points (2, 9), and (5, 12)? (1 point)

OpenStudy (anonymous):

1is the ans

OpenStudy (anonymous):

12 - 9 = 3 5 - 2 = 3 3/3 = 1

OpenStudy (anonymous):

\[\frac{ y_2 - y_1 }{ x_2 - x_1 } = \frac{ 12-9 }{ 5-2 } = \frac{ 3 }{ 3 } = 1\]